# Orion: A Fully Homomorphic Encryption Framework for Deep Learning

### v2

@@ 4.3. Single-Shot Multiplexed Convolutions
We solve this problem by recognizing that the rows of any Toeplitz matrix can be permuted without changing its underlying computation. This lets us relax the invariant that our ciphertext must represent the image as a flattened tensor. Our approach builds on the multiplexed parallel convolutions introduced by Lee et al. (lee2022, ), which interleaves channels within a ciphertext to efficiently handle strided convolutions. Notably, their method consumes two multiplicative levels: one level to first perform a non-strided convolution, and a second level to mask and collect only the correct elements (see Figure 5 of (lee2022, )). On the other hand, our single-shot multiplexed packing strategy consumes just one level by fusing this mask-and-collect step directly into the weight matrix, which can be pre-processed.
Figure 4b shows our method for converting the standard Toeplitz matrix in Figure 4a to our single-shot multiplexed solution. Our method produces a densely packed output ciphertext, parameterized by a gap, $g$. Subsequent non-strided convolutions maintain this gap, while strided convolutions increase it by a factor of $s$. This reformulation lets us leverage both BSGS and double-hoisting and doing so reduces the number of ciphertext rotations in ResNet-20 (he2015deepresiduallearningimage, ) from $1457$ to $836$. Table 2 of our results further compares these two packing strategies in larger neural networks.
[…]
@@
5. Automatic Bootstrap Placement
In this section, we discuss the problem of bootstrap placement in private neural inference. We begin by highlighting its barrier to entry to motivate the need for automated solutions. Then, we describe the challenges that arise in automation. We close with a general and efficient solution for modern neural networks based on topological sorting.
5.1. Problem Overview
Recall from Section 2 that the goal of CKKS bootstrapping is to enable further homomorphic multiplications by increasing a ciphertext’s level from $\ell=0$ to $\ell=L_{\text{eff}}<L$. Determining the optimal location of bootstrap operations within a network is perhaps the most challenging aspect of FHE inference for three reasons. First, placing a bootstrap at any point in the network directly affects the levels of all subsequent operations, including future bootstrap locations. Second, naïve strategies that delay bootstrapping until absolutely necessary often result in more bootstraps being placed, particularly in networks with residual connections (see Figure 10 of Fhelipe (fhelipe, )). And finally, bootstrap runtime grows superlinearly with $L_{\text{eff}}$. Thus, only minimizing the number of bootstrap operations often counter-intuitively increases network latency.
In Orion, our goal is to automatically determine the locations of bootstrap operations to minimize network latency. In doing so, we also determine the levels of all linear layers and activation functions. Collectively, we refer to our decisions as a level management policy.
Our constraints: To efficiently realize this goal, we impose three constraints. First, we restrict the placement of bootstrap operations between network layers. This greatly simplifies our analysis without significantly impacting its quality, since optimal policies inherently avoid bootstrapping within layers, where circuits are the widest. Second, we neglect the cost of activation functions and only model the latencies of evaluating linear layers and bootstraps. Indeed, $92.6$% of our ResNet-20 latency comes from these two sources. Finally, we exclude networks with overlapping skip connections from our analysis, such as DenseNets (densenet, ).
@@ 5.2. Orion’s Level Management Policy
The core idea: To introduce our level management policy, consider the simple three-layer fully connected network in Figure 5 without intermediate activation functions. The linear layers $(\texttt{fc1-}\texttt{fc3})$ each consume one level, and we set the effective level ($L_{\text{eff}}$) to $3$. By restricting the placement of bootstrap operations to between network layers, we make it possible to efficiently construct a directed acyclic graph (DAG) that enumerates all possible network states. We refer to these graphs as level digraphs, and the level digraph for the network in Figure 5 is shown in Figure 5. Each row of nodes in Figure 5 represents one linear layer. Nodes within a row are each assigned a valid level and weighted by the latency of performing their corresponding linear layer at that level. A valid level $\ell$ satisfies $d\leq\ell\leq L_{\text{eff}}$, where $d$ is the layer’s multiplicative depth. Edges between nodes are weighted by the latency of bootstrapping up to $L_{\text{eff}}$ if it is required. In Figure 5, we highlight the edges that require bootstrap operations in red.
@@
A general strategy: One problem with the approach in Figure 5 is that it does not support networks with multiple paths from input to output. Consider the network in Figure 5 with a residual connection and an activation function. Residual connections are a staple of modern neural network design, however complicate the choice of level management policy. A viable solution to Figure 5 might include solving two independent level digraphs, one per path through the network. However, this approach can lead to clashing choices for the levels of nodes common to both paths.
[…]
Implementation details: We estimate the latencies of both the linear layers and bootstrap operations with an analytical model. Then, Orion automatically inserts bootstrap operations within the network, abstracting these complexities away from the end-user. Orion’s automatic level management and bootstrap placement takes only $6.67$ ($49.4$) seconds for ResNet-20 (ResNet-110) and scalability is further explored in Table 4 of our results.
@@ 6. The Orion Framework
We now present the Orion framework and API, which includes automated support for the optimizations and techniques presented above (i.e., bootstrap placement and packing strategies). In addition, Orion automatically handles several subtle, yet difficult, challenges that arise when building large-scale FHE programs: range estimation for high-precision bootstrapping and Chebyshev polynomial evaluation, error-free scaling factor management, and system support for very large data structures.
@@
The Orion API: FHE neural network applications are notoriously difficult to program, and significant research efforts  (porcupine, ; coyote, ; CHET, ; EVA, ; viand2023heco, ; fhelipe, ) have focused on reducing this barrier to entry. Yet, despite this progress, no high-level, high-performance API exists, which practitioners have come to expect from modern tools like JAX or Pytorch (jax2018github, ; pytorch, ). Orion provides this interface. We implement custom Python bindings to the Lattigo (lattigo, ) FHE library that directly interact with PyTorch tensors. Then, we use these bindings to build Orion modules that inherit and extend the functionality of their corresponding PyTorch modules. For example, Listing 1 shows a ResNet block in Orion that closely mirrors the equivalent PyTorch implementation. Orion facilitates PyTorch’s ease of use while achieving state-of-the-art FHE neural network performance. While we chose PyTorch, it is straightforward to add support for other modern deep learning libraries.
[…]
[python] fontsize=, fontfamily=lmtt, escapeinside=——,
{minted} [linenos=false]python —import— —torch.nn— —as— —nn— —import— —orion.nn— —as— —on
class— —BasicBlock—(on.Module): —def— —__init__—(—self—, Ci, Co, stride=1): —super—().—__init__—() —self——.conv1 = on.Conv2d(Ci, Co, 3, stride, 1)— —self—.bn1 = on.BatchNorm2d(Co) —self—.act1 = on.ReLU(degrees=[15,15,27])
self—.conv2 = on.Conv2d(Co, Co, 3, 1, 1) —self—.bn2 = on.BatchNorm2d(Co) —self—.act2 = on.SiLU(degree=63)
self—.add = on.Add() —self—.shortcut = nn.Sequential() —if— stride != 1: —self—.shortcut = nn.Sequential( on.Conv2d(Ci, Co, 1, stride, 0), on.BatchNorm2d(Co))
def— —forward—(—self—, x): out = —self—.act1(—self—.bn1(—self—.conv1(x))) out = —self—.bn2(—self—.conv2(out)) out = —self—.add(out, —self—.shortcut(x)) —return— —self—.act2(x)
Listing 1 ResNet block instantiation in Orion.
@@
Range estimation: Both high-precision bootstrapping and high-degree (Chebyshev) polynomial activation functions require the inputs to be in the range $[-1,1]$. Thus, when building large FHE programs, practitioners must insert specific scale-down operations to ensure this property holds for all calls to bootstrapping and Chebyshev evaluations. This step is especially necessary for neural networks, which generate intermediate activation values outside of $[-1,1]$. Prior work chooses per network scaling factors by manually inspecting and tracking the largest intermediate values seen (lee2022, ; Kim2023, ). Orion handles this process automatically through net.fit(), which accepts the entire training dataset as input, calculates per layer scaling factors, and inserts scale-down multiplications directly into the computational graph.
@@
Scale management: Recall from Section 2.2 that encoding a cleartext into a CKKS plaintext involves multiplication by a scaling factor, $\Delta$. Properly managing this scale factor for large FHE programs is challenging. To highlight this, consider performing a homomorphic multiplication (either $\mathsf{PMult}$ or $\mathsf{HMult}$) where both operands are at level $\ell$ and have a scaling factor of $\Delta$. The resulting ciphertext has a scaling factor of $\Delta^{2}$ that is then rescaled to $\Delta^{2}/q_{\ell}\approx\Delta$ where $q_{\ell}$ is the last prime limb. Thus, rescaling also introduces approximation errors into the resulting ciphertext which grows with circuit depth (errorreduced, ).
@@
In Orion, we propose a new technique to automatically handle scale management that we call errorless neural network evaluation. At a high level, we maintain the invariant that the scaling factor of any ciphertext between network layers must be precisely $\Delta$. This technique extends the errorless, depth-optimal polynomial evaluation (DHBSGS, ) to full neural network inferences. As an example of our technique, assume we are performing a convolutional in Orion. We leverage the fact that our compilation phase has already pre-determined the level at which to perform this convolution (we will call this level $j$). We choose to then encode its weights with scale factor $q_{j}$ (rather than $\Delta$), the last RNS modulus at level $j$. Performing this convolution with a ciphertext at scale $\Delta$ results in an output with scale $\Delta\cdot q_{j}$. Rescaling will divide the ciphertext by the last prime modulus, $q_{j}$, thus precisely resetting this output ciphertext scale back to $\Delta$.
@@
Handling large data structures: Large datasets and networks require hundreds of gigabytes of rotation keys and matrix diagonals. Orion provides support to store these large data structures to disk in HDF5 format (hdf5, ). Then, the correct data is loaded dynamically during inference to minimize the size of our transient data.
@@ 7. Methodology
CPU setup: We evaluate Orion on a C4 GCP instance with an Intel Xeon Platinum $8581$C processor clocked at $2.3$ GHz and $512$ GB of RAM. We target Lattigo v5.0.25.0.25.0.25.0.2 (lattigo, ) in our evaluation, an open-source, single-threaded FHE library that supports both double-hoisted matrix-vector products and errorless, depth-optimal polynomial evaluation. Future work includes supporting multi-threaded FHE backends (e.g., OpenFHE (OpenFHE, )) and GPU implementations (e.g., HEAAN (cheon2017homomorphic, )). For a fair comparison, we rerun Lee et al. (lee2022, ) and Fhelipe (fhelipe, ) on the exact same C4 GCP instance, which are also single-threaded.
@@
Activation functions: As shown in Table 1, we use the $x^{2}$ activation function for MNIST networks and either ReLU or its smoother variant, SiLU (silu, ) for all other datasets and networks. We approximate ReLU via a minimax composite polynomial (lee2022, ). Here, a high degree polynomial is decomposed into a composition of several lower degree polynomials to reduce the number of homomorphic multiplications. Following prior work, our composition consists of three polynomials of degree $15$, $15$, and $27$. For SiLU, we use a degree-$127$ Chebyshev polynomial, obtained using a similar minimax approach. The multiplicative depth of ReLU is $14$ ($13$ for $\mathsf{sign}$($x$) and $1$ for its multiplication with $x$), whereas the depth of SiLU is just $7$. Later, we will show how the choice of activation function introduces a trade-off in latency and test accuracy.
@@
Benchmarks: Our evaluation consists of four datasets (image sizes): MNIST ($28\times 28\times 1$) (lecun-mnisthandwrittendigit-2010, ), CIFAR-10 ($32\times 32\times 3$) (krizhevsky2009learning, ), Tiny ImageNet ($64\times 64\times 3$) (Le2015TinyIV, ), and ImageNet ($224\times 224\times 3$) (imagenet, ). We evaluate three neural networks for MNIST from prior work: a 3-layer MLP from SecureML (secureML, ), a 3-layer CNN from LoLA CryptoNets (Brutzkus2019LowLatency, ), and the largest LeNet-5 model from CHET (CHET, ) and EVA (EVA, ). For CIFAR-10, we benchmark AlexNet (alexnet, ), VGG-16 (vgg, ), and ResNet-$20$ (he2015deepresiduallearningimage, ), each with ReLU and SiLU activation functions. For Tiny ImageNet, we implement MobileNet-v1 (mobilenet, ) and ResNet-$18$, and for ImageNet, we evaluate ResNet-$34$ and ResNet-$50$. For all networks, we replace max pooling with average pooling.
@@
Validation: We validate each benchmark by comparing Orion’s FHE outputs against the equivalent cleartext outputs from PyTorch. For MNIST, we perform this validation across all 10,000 test images. For CIFAR-10 (Tiny ImageNet), we instead randomly sample 1,000 ($100$) test images. And for ImageNet, we perform just a single encrypted inference for comparison due to resource constraints.
@@
Alongside accuracy and latency, we also report the mean precision (in bits) of the output, which is defined as $-\log_{2}(\epsilon)$, where $\epsilon$ is the mean absolute difference between the outputs of Orion and PyTorch. To compare with prior work, we also report the number of ciphertext rotations and number of bootstrap operations for each network. The links to code blocks for each network and their respective parameter sets can be found in Table 1.
@@
MLP $0.12$ $0.12$ $70$ $x^{2}$ $5$ 00 $98.02\%$ $98.03\%$ $4.60$ $0.29$
LoLA $0.10$ $0.13$ $73$ $x^{2}$ $5$ 00 $98.63\%$ $98.62\%$ $4.81$ $0.23$
MNIST LeNet $1.66$ $4.30$ $282$ $x^{2}$ $7$ 00 $99.31\%$ $99.31\%$ $10.4$ $2.93$

### v3

@@ 4.3. Single-Shot Multiplexed Convolutions
We solve this problem by recognizing that the rows of any Toeplitz matrix can be permuted without changing its underlying computation. This lets us relax the invariant that our ciphertext must represent the image as a flattened tensor. Our approach builds on the multiplexed parallel convolutions introduced by prior work (Lee et al., 2022b), which interleaves channels within a ciphertext to efficiently handle strided convolutions. Notably, their method consumes two multiplicative levels: one level to first perform a non-strided convolution and a second level to mask and collect only the correct elements (see Figure 5 of (Lee et al., 2022b)). On the other hand, our single-shot multiplexed packing strategy consumes just one level by fusing this mask-and-collect step directly into the weight matrix, which can be pre-processed.
Figure 5b shows our method for converting the standard Toeplitz matrix in Figure 5a to our single-shot multiplexed solution. Our method produces a densely packed output ciphertext, parameterized by a gap, $g$. Subsequent non-strided convolutions maintain this gap, while strided convolutions increase it by a factor of $s$. This reformulation lets us leverage both BSGS and double-hoisting and doing so reduces the number of ciphertext rotations in ResNet-20 (He et al., 2015) from $1457$ to $836$. Table 3 of our results further compares these two packing strategies in larger neural networks.
Multi-ciphertext: When an input image does not fit into a single ciphertext (i.e., the number of pixels is greater than the number of slots), the image is split across multiple ciphertexts into contiguous slots where the last ciphertext may only be partially filled with data. In this case, we naturally perform a blocked matrix-vector product with blocks of size $\textit{slots}\times\textit{slots}$. While we evaluate Orion in a single-threaded setting, each block performs independent work and is well-suited for parallel execution across multiple threads or cores.
5. Automatic Bootstrap Placement
In this section, we discuss the problem of bootstrap placement in private neural inference. We begin by highlighting the challenges to motivate the need for automated solutions. Then, we describe the challenges that arise in automation. We close with a general and efficient solution for modern neural networks based on solving shortest paths in directed acyclic graphs (DAGs).
5.1. Problem Overview
Recall from Section 2 that the goal of CKKS bootstrapping is to enable further homomorphic multiplications by increasing a ciphertext’s level from $\ell=0$ to $\ell=L_{\text{eff}}<L$. Determining the optimal location of bootstrap operations within a network is a challenging aspect of FHE inference for three reasons. First, placing a bootstrap at any point in the network directly affects the levels of all subsequent operations, including future bootstrap locations. Second, naïve strategies that delay bootstrapping until absolutely necessary often result in more bootstraps being placed, particularly in networks with residual connections (e.g., see Figure 10 of Fhelipe (Krastev et al., 2024)). Finally, bootstrap runtime grows superlinearly with $L_{\text{eff}}$, as shown in Figure 1. Thus, only minimizing the number of bootstrap operations often counter-intuitively increases network latency, since it increases per-bootstrap costs dramatically and increases the average level (i.e., latency) of all homomorphic operations, as shown in Figures 1 and 1. Therefore, a careful balance must be struck to have a high enough $L_{\text{eff}}$ in order to perform meaningful computation without incurring an exorbitant cost of bootstrapping.
In Orion, our goal is to automatically determine the locations of bootstrap operations to minimize network latency. In doing so, we also determine the levels to perform all linear layers and activation functions. Collectively, we refer to our decisions as a level management policy.
@@ 5.1. Problem Overview
Placement constraint: We restrict the placement of bootstrap operations to occur between network layers, where we define a network layer to be a linear transformation or a polynomial evaluation. Since all linear transforms (e.g., convolutions and fully-connected layers) in Orion consume one level, there is no bootstrap decision to be made within them. On the other hand, activation functions in CKKS must be expressed as a polynomial (e.g., SiLU) or as a composition of several polynomials (e.g., ReLU, which consists of three polynomials (Lee et al., 2022b)) and these approximations tend to consume many levels. We restrict Orion to bootstrap between but not within polynomial evaluations.
At a high level, a CKKS polynomial evaluation takes as input a single ciphertext and applies the polynomial element-wise to each slot. The evaluation fans out into a wider circuit of intermediates during the reduction process, which outputs again a single ciphertext. Therefore, if we did choose to bootstrap any of these intermediate ciphertexts, all intermediates at that level would also have to be bootstrapped, increasing bootstrap calls. The reduction steps of polynomial evaluation can be found in prior work (Bossuat et al., 2020); we refer the interested reader to Line 4 of Algorithm 1 and Lines $7$, $15$, $17$, 1919 of Algorithm $2$. Thus, we do not bootstrap within polynomials, which also simplies our analysis. Finally, we exclude networks with overlapping skip connections from our analysis, such as DenseNets (Huang et al., 2017).
Cost model: Our cost model accounts for the depth of activation layers, the depth and latency of all linear layers, and bootstrap latency. This cost model captures the depth of the circuit (a majority of which comes from the activations) as well as the latency (stemming from the linear layers and bootstrapping). Empirically, we find that 94.194.1% (95.795.7%) of our ResNet-20 (ResNet-18) latency comes from these two sources.
@@ 5.2. Orion’s Level Management Policy
The core idea: To introduce our level management policy, consider the simple three-layer fully connected network in Figure 6 without intermediate activation functions. The linear layers $(\texttt{fc1-}\texttt{fc3})$ each consume one level, and we set the effective level ($L_{\text{eff}}$) to $3$. By restricting the placement of bootstrap operations to between network layers, we make it possible to efficiently construct a directed acyclic graph (DAG) that enumerates all possible network states. We refer to these graphs as level digraphs, and the level digraph for the network in Figure 6 is shown in Figure 6. Each row of nodes in Figure 6 corresponds to one linear layer. Nodes within a row represent possible choices of level for that linear layer and are weighted by the latency of performing their corresponding linear layer at that level (recall from Figure 1 that the latencies of homomorphic operations vary as a function of $\ell$). A valid level $\ell$ satisfies $d\leq\ell\leq L_{\text{eff}}$, where $d$ is the layer’s multiplicative depth. The edges between nodes are weighted by the latency of bootstrapping up to $L_{\text{eff}}$ if required. In Figure 6, we highlight the edges that require bootstrap operations in red.
Note that even when a bootstrap occurs, the subsequent layer can still be performed at $\ell<L_{\text{eff}}$. While this choice may seem wasteful, the additional degree of freedom often leads to more efficient level management policies, especially in more realistic networks (e.g., ResNet-20). The optimal level management policy in Figure 6 contains the operations along the shortest (lowest total weight) path from any input node ($\texttt{input}_{0\leq\ell\leq 3}$) to any output node ($\texttt{fc3}_{1\leq\ell\leq 3}$) in the level digraph, minimizing the latency (with respect to our heuristics) of an inference.
A general strategy: One problem with the approach in Figure 6 is that it does not support networks with multiple paths from input to output. Consider the network in Figure 6 with a residual connection and an activation function. Residual connections are a staple of modern neural network design but complicate the choice of level management policy. A viable solution to Figure 6 might include solving two independent level digraphs, one per path through the network. However, this approach can lead to clashing choices for the levels of nodes common to both paths.
Instead, we observe that each residual connection forms a single-entry, single-exit (SESE) region (Johnson et al., 1994) in the network graph, bounded by pairs of fork and join nodes. Here, a fork (join) node is any node with more than one child (parent). In Figure 6, the nodes {input} and {+} are the only such fork and join nodes. We leverage this property to extract any SESE region around a residual connection (including its fork and join nodes), treat it as its own distinct problem, and then black-box its solution. To do this, we first construct two separate level digraphs: one for the backbone network and one for the residual connection. Then, we collapse these separate digraphs into one aggregate level digraph by solving a joint shortest path problem between every pair of input and output nodes. For instance, in Figure 6, the weight of the edge in the ”aggregate” digraph from $\texttt{\{input\}}_{\ell=0}$ to $\texttt{\{+\}}_{\ell=0}$ is set to the sum of (i) the total weight of the shortest path from $\texttt{\{input\}}_{\ell=0}$ to $\texttt{\{+\}}_{\ell=0}$ within the ”backbone” digraph and (ii) the total weight of the shortest path from $\texttt{\{input\}}_{\ell=0}$ to $\texttt{\{+\}}_{\ell=0}$ within the ”residual” digraph.
Finally, we insert this aggregate level digraph back into the larger network, allowing us to find a shortest path similarly to Figure 6. By also tracking the paths summarized by each edge in every aggregate level digraph, we can then work backwards to determine the correct level management policy for all layers, including within the SESE regions that we black-boxed.
Implementation details: In CKKS, the ReLU activation function is approximated by $x\cdot\mathsf{sign}(x)$, which is itself a SESE region. As ReLUs are often placed within larger residual connections, our approach always begins by black-boxing the inner-most SESE regions before working outwards.
We estimate the latencies of both the linear layers and bootstrap operations with an analytical model. Then, Orion automatically inserts bootstrap operations within the network, abstracting these complexities away from the end-user. Orion’s automatic level management and bootstrap placement algorithm takes only $1.94$ ($0.82$) seconds for ResNet-20 (AlexNet), which is an $8.14\times$ ($1270\times$) speed-up over Dacapo (Cheon et al., 2024) on these same networks. To demonstrate that our solution is tractable, we also ran the algorithm on ResNet-1202, which took just $151$ seconds. Scalability with network depth is further explored in Table 5 of our results.
@@
Listing 1 ResNet block instantiation in Orion.
@@ 6. The Orion Framework
We now present the Orion framework and API, which includes automated support for the optimizations and techniques presented above (e.g., bootstrap placement and packing strategies). In addition, Orion automatically handles several subtle, yet difficult, challenges that arise when building large-scale FHE programs: range estimation for high-precision bootstrapping and Chebyshev polynomial evaluation, error-free scaling factor management, and system support for very large data structures.
@@
The Orion API: FHE neural network applications are notoriously difficult to program, and significant research efforts  (Cowan et al., 2021; Malik et al., 2023; Dathathri et al., 2019, 2020; Viand et al., 2023; Krastev et al., 2024) have focused on reducing this barrier to entry. Yet, despite this progress, no high-level, high-performance API exists, which practitioners have come to expect from modern tools like JAX or Pytorch (Bradbury et al., 2018; Paszke et al., 2019). Orion provides this interface. We implement custom Python bindings to the Lattigo (lat, 2022) FHE library that directly interact with PyTorch tensors. Then, we use these bindings to build Orion modules that inherit and extend the functionality of their corresponding PyTorch modules. For example, Listing 1 shows a ResNet block in Orion that closely mirrors the equivalent PyTorch implementation (  ). In more detail, we make use of our custom orion.nn module (rather than torch.nn) when instantiating specific layers. Within orion.nn, activation functions accept a keyword argument that specifies the degree of each composite polynomial used in the approximation. Orion facilitates PyTorch’s ease of use while achieving state-of-the-art FHE neural network performance. While we chose PyTorch, it is straightforward to add support for other modern deep learning libraries.
@@
Range estimation: Both high-precision bootstrapping and high-degree (Chebyshev) polynomial activation functions require the inputs to be in the range [1,1][-1,1]. Thus, when building large FHE programs, practitioners must insert specific scale-down $\mathsf{PMults}$ to ensure this property holds for all calls to bootstrapping and Chebyshev evaluations. This step is especially necessary for neural networks, which generate intermediate activation values outside of [1,1][-1,1].
Prior work chooses per network scaling factors by manually inspecting and tracking the largest intermediate values seen (Lee et al., 2022b; Kim and Guyot, 2023). Orion handles this process automatically through net.fit(), which accepts the entire training dataset as input, calculates per layer scaling factors, and inserts scale-down multiplications directly into the computational graph. Then, all standard (e.g., ReLU) element-wise activation functions are fit with Chebyshev polynomials, either through interpolation or by the Remez algorithm (Lee et al., 2022c). Listing 1 shows Orion’s support for both ReLU and SiLU where the degree(s) of their approximation are passed in as keyword arguments. Extending support to other activation functions is straightforward and follows a process similar to defining custom PyTorch modules. The user need only extend the base orion.nn module, inheriting support for range estimation and polynomial evaluation, and provide an activation function to approximate with a specified degree. In this way, Orion is able to support arbitrary activation functions that can be fit with high-degree polynomials.
Figure 7. An example of Orion’s scale management policy. By encoding plaintexts ($\mathbf{w}]$) using the appropriate $q_{j}$ that will be used during rescaling, we are able to maintain the invariant that the scale of any ciphertext must be precisely $\Delta$ between network layers. Each edge color in the graph corresponds to a CKKS polynomial at a specific scale value.
@@
Scale management: Recall from Section 2.2 that encoding a cleartext into a CKKS plaintext involves multiplication by a scaling factor, $\Delta$. Properly managing this scale factor for large FHE programs is challenging. To highlight this issue, consider performing a homomorphic multiplication (either $\mathsf{PMult}$ or $\mathsf{HMult}$) where both operands are at level $\ell$ and have a scaling factor of $\Delta$. The resulting ciphertext has a scaling factor of $\Delta^{2}$ that is then rescaled to $\Delta^{2}/q_{\ell}\approx\Delta$ where $q_{\ell}$ is the last prime limb. Thus, rescaling also introduces approximation errors into the resulting ciphertext which grows with circuit depth (Kim et al., 2020).
@@
In Orion, we propose a new technique to automatically handle scale management that we call errorless neural network evaluation. At a high level, we maintain the invariant that the scaling factor of any ciphertext between network layers must be precisely $\Delta$. This technique extends the errorless, depth-optimal polynomial evaluation from Bossuat et al. (Bossuat et al., 2020) to full neural network inferences. Figure 7 shows an example of our technique: assume we are performing a convolution in Orion. We leverage the fact that our compilation phase has already pre-determined the level at which to perform this convolution (we will call this level $j$). We choose to then encode the kernel weights with scale factor $q_{j}$ (rather than $\Delta$), the last RNS modulus at level $j$. Performing this convolution with a ciphertext at scale $\Delta$ results in an output with scale $\Delta\cdot q_{j}$. Rescaling will divide the ciphertext by the last prime modulus, $q_{j}$, thus precisely resetting this output ciphertext scale back to $\Delta$.
@@
Handling large data structures: Large datasets and networks require hundreds of gigabytes of rotation keys and matrix diagonals. Orion provides support to store these large data structures to disk in HDF5 format (Folk et al., 2011). Then, the correct data can be loaded dynamically during inference to minimize the size of our transient data.
@@ 7. Methodology
CPU setup: We evaluate Orion on a C4 GCP instance with an Intel Xeon Platinum 85818581C processor clocked at $2.3$ GHz and $512$ GB of RAM. We target Lattigo v5.0.25.0.2 (lat, 2022) in our evaluation, an open-source, single-threaded FHE library that supports both double-hoisted matrix-vector products and errorless, depth-optimal polynomial evaluation. Future work includes supporting multi-threaded FHE backends (e.g., OpenFHE (Badawi et al., 2022)) and GPU implementations (e.g., HEaaN (HEA, 2016)). For a fair comparison, we rerun Lee et al. (Lee et al., 2022b) and Fhelipe (Krastev et al., 2024) on the exact same C4 GCP instance, which are also single-threaded.
@@
Activation functions: As shown in Table 2, we use the $x^{2}$ activation function for MNIST networks and either ReLU or its smoother variant, SiLU (Elfwing et al., 2017), for all other datasets and networks. We approximate ReLU via a minimax composite polynomial (Lee et al., 2022c). Here, a high degree polynomial is decomposed into a composition of several lower degree polynomials to reduce the number of homomorphic multiplications. Following prior work, our composition consists of three polynomials of degree $15$, $15$, and $27$. For SiLU, we use a degree-$127$ Chebyshev polynomial, obtained using a similar minimax approach. The multiplicative depth of ReLU is $14$ ($13$ for $\mathsf{sign}$($x$) and $1$ for its multiplication with $x$), whereas the depth of SiLU is just $7$. Later, we will show how the choice of activation function introduces a trade-off in latency and test accuracy.
@@
Benchmarks: Our evaluation consists of four datasets (image sizes): MNIST ($28\times 28\times 1$) (LeCun and Cortes, 2010), CIFAR-10 ($32\times 32\times 3$) (Krizhevsky et al., 2009), Tiny ImageNet ($64\times 64\times 3$) (Le and Yang, 2015), and ImageNet ($224\times 224\times 3$) (Deng et al., 2009). We evaluate three neural networks for MNIST from prior work: a 3-layer MLP from SecureML (Mohassel and Zhang, 2017), a 3-layer CNN from LoLA CryptoNets (Brutzkus et al., 2019), and the largest LeNet-5 model from CHET (Dathathri et al., 2019) and EVA (Dathathri et al., 2020). For CIFAR-10, we benchmark AlexNet (Krizhevsky et al., 2012), VGG-16 (Simonyan and Zisserman, 2015), and ResNet-$20$ (He et al., 2015), each with ReLU and SiLU activation functions. For Tiny ImageNet, we implement MobileNet-v1 (Howard et al., 2017) and ResNet-$18$, and for ImageNet, we evaluate ResNet-$34$ and ResNet-$50$. For all networks, we replace max pooling with average pooling. Finally, we perform object detection and localization using YOLO-v1 (Redmon et al., 2016) with a ResNet-34 backbone on the PASCAL-VOC dataset ($448\times 448\times 3$) (Everingham et al., 2010).
@@
Validation: We validate each benchmark by comparing Orion’s FHE outputs against the equivalent cleartext outputs from PyTorch. For MNIST, we perform this validation across all 10,000 test images. For CIFAR-10 (Tiny ImageNet), we instead randomly sample 1,000 ($100$) test images, leading to an estimate of the test accuracy. Finally, for ImageNet, we perform just a single encrypted inference for comparison due to resource constraints.
@@
Alongside accuracy and latency, we also report the mean precision (in bits) of the output, which is defined as $\hskip-1.00006pt\log_{2}(\epsilon)$, where $\epsilon$ is the mean absolute difference between the outputs of Orion and PyTorch. To compare with prior work, we also report the number of ciphertext rotations and number of bootstrap operations for each network. The links to code blocks for each network and their respective parameter sets can be found in Table 2.
@@
MLP $0.12$ $0.12$ $70$ $x^{2}$ 55 0 $98.02\%$ $98.03\%$ $4.60$ $0.29$
LoLA $0.10$ $0.13$ $73$ $x^{2}$ 55 0 $98.63\%$ $98.62\%$ $4.81$ $0.23$
MNIST LeNet $1.66$ $4.30$ $282$ $x^{2}$ 77 0 $\% 99.3199.31\%$ $10.4$ $2.93$